# Physical properties of strong 1 ¡ z ¡ 3 Balmer and Paschen lines emitters observed with JWST

### v1

@@ 6.2 Derivation of the effective attenuation curves
With CIGALE, we can derive the effective attenuation curve for each galaxy of our subsample, being the total attenuation in the V band. We selected fifteen wavelength bands (from the rest-frame far-UV to the near-IR) as the code is able to compute the attenuation in filters different from the ones used to fit the input data. With this feature in mind we added fifteen custom-made filters to the code. We choose to create very narrow filters (10 nm each) to carefully select which part of the spectrum is going to be sampled. The reason for this choice is two-fold: we can avoid contamination from emission and absorption lines but also pinpoint a specific part of the spectrum. We use the bayesian estimations of CIGALE for fifteen photometric bands to reconstruct the total effective attenuation curve of our subsample of galaxies.
@@
and we report the resulting exponents in Table 6. We also fit our whole subsample with a single powerlaw and find an exponent of =-0.96 0.10. By colour-coding each curve with its associated value of , we see a trend between the decrease in and the steepening of the curve (i.e. the exponent is closer to -1.3). A high equals a negligible increase of attenuation for stars in birth clouds compared to the ISM meaning a single powerlaw (for ID 8588). The differential attenuation found for the galaxies with ¡ 0.35 mag leads to a steeper attenuation law. Using radiative transfer modelling, it has been shown that the effective attenuation curve depends on the amount of dust attenuation (e.g. Pierini et al., 2004; Saftly et al., 2015; Feldmann et al., 2017; Roebuck et al., 2019; Trayford et al., 2020). As predicted by the models, the IR part of the curves do not follow the model of Calzetti et al. (2000) and ID 8588 is the only galaxy to follow the model in the UV-optical which shows the need for a two-components dust attenuation model. Finally, we compared the values of obtained through a fit made with , , and being free and a fit made with and being free and both exponents fixed to -0.96. Both sets of values are consistent within their respective errorbars.
@@ 7 Impact of input data type and models on SFRs and stellar mass determinations
As getting robust SFRs and stellar masses is crucial in numerous contexts (e.g. Pacucci et al., 2023; Juodžbalis et al., 2024), we explore the effects of the choice of stellar populations to fit the spectro-photometric data on both aforementioned output parameters. Then, we quantify the variation of SFRs and stellar masses between a fit performed with broad-band photometry alone and with spectro-photometric data, the latter being used as our reference fit. Finally, we assess the impacts of the change of the attenuation recipe on the SFRs and stellar masses when fitting spectro-photometric data with the BC03 models.
@@ 7.1 Fits with BPASS or BC03 models
Here we compare SFRs and stellar mass estimations obtained with either BPASS or BC03 models. We start with SFRs estimations. Fig 12 shows a clear trend of increased SFRs when using BPASS compared to BC03 models. The SFRs derived through SED fitting with BPASS are, on average, 0.34 dex larger than with BC03 models. We attribute this increase in SFR to the higher emission line intensity of the BPASS models addressed in Sect 4.1. We searched for a correlation between the variation of SFRs and other output parameters and found a fraction of burst 5 times higher with BPASS for the most deviant object (ID 10293, red dot) in Fig 12.
We find higher stellar masses with BC03 models as shown in Fig 12. The mass derived through SED fitting with the BPASS models is, on average, 0.18 dex lower than with the BC03 models.Osborne & Salim (2024) also report lower masses (by 0.17 dex) when using the BPASS models compared to the BC03 models for a sample of CANDELS sources with redshift between 0.7 and 2.3. We attribute these lower masses to the deficit in low mass stars of the BPASS models addressed in Sect 4.2. We also look for correlations between the change in stellar mass and other parameters. The SFH of the galaxy exhibiting the largest mass difference differs a lot between the runs with BC03 and BPASS as ID 10293 experiences a strong change in burstiness as mentioned in the previous paragraph.
@@ 7.2 Fits of photometric data or spectro-photometric data
Here we assess the impact of using only photometry with both stellar populations models on the stellar mass and SFR estimations. In Fig 13, we compare the estimates of the SFR obtained with a fit of photometric data to those from a fit of spectro-photometric data with both BC03 and BPASS. The SFR derived with the fit of photometric data only is, on average, 0.16 dex higher with BC03 and 0.03 dex higher with BPASS compared to values obtained with spectro-photometric measurements. Considering the scatter found in this comparison we also provide the maximal and minimal absolute value of the difference in . For BC03 (BPASS), the largest discrepancy is of 0.42 (0.60) dex and the lowest is of 0.04 (0.09) dex. Errors on the SFRs also become much lower when fitting the spectro-photometric data. SFRs lower than 30 M⊙/yr tend to be systematically overestimated when only photometric fluxes are fitted by 0.18 dex with BC03 and by 0.09 dex with BPASS. We attribute this increase to the stronger emission lines computed by the code in the absence of spectroscopic data as shown in Fig 9 and which are due to a more active star formation process. Our results are in agreement with Pacifici et al. (2015) who also found a higher SFR (by 0.13 dex) while using BC03 and photometric data only for a sample of 364 3D-HST galaxies at z ¡ 3.
The stellar masses are found larger when EWs are added to the fit by 0.15 dex regardless of the chosen stellar populations as shown in Fig 13. The lower masses obtained with only photometric data are due to an overestimation of the line contribution (see Fig 9): the contribution of young stars to the SED is higher leading to a lower contribution of older stars which make the bulk of the galaxy mass. Our results are in agreement with the findings of Schaerer & de Barros (2010) who fitted z 7 sources while Pacifici et al. (2015) found no noticeable difference in the stellar mass output with the addition of spectroscopic data to the fits for their sample of z ¡ 3 galaxies.
@@ 7.3 Fits with different attenuation recipes
Dust attenuation has a direct impact on the derived SFRs and, to a lesser extent, on stellar masses. Any change in the shape of the attenuation curve can affect (albeit in varying degrees) the estimations of these quantities as shown in (e.g. Schaerer & de Barros, 2010; Pacifici et al., 2015; Lo Faro et al., 2017; Buat et al., 2018, 2019). In the following analysis, we compare the effects of changing the attenuation curve parameters in the framework of the CF00 recipe.
We test three different attenuation curves with separate runs using the BC03 models only and the full spctro-photometric dataset. We do not include the BPASS models in these tests, as the attenuation curves presented in Sect. 6.2 cannot be reconstructed with the binary stellar populations models. The first run, called , is performed with the four parameters of the modified CF00 recipe taken free as presented in Table 4; it will be considered the reference run. The second run is performed with the original CF00 recipe () in which is fixed to 0.3 and both exponents are equal to -0.7 while the third one makes use of the single average power-law introduced in Sect. 6.2, is taken free and we call this run .
As already noted in Sect. 6.2, we do not find noticeable changes in between and . The SFRs are found consistent with a slightly larger scatter with CF00 than with the single power-law as shown in Fig 14. The agreement between the 3 stellar mass estimations is very good with a scatter of 0.03 and 0.06 dex for and respectively, when they are compared to (see Fig 14).
From the results shown in this section, we conclude that fixing the slope of the attenuation curve (and even ) is much less impactful on the SFRs and stellar masses than including binary stellar populations or fitting only photometric data. However, this lesser effect could be a result of our sample being mostly composed of objects with low dust content. Indeed, the most discrepant objects in Fig 14 have an amount of dust ¿ 1 mag.
@@ 8 Conclusions
In this work, we assessed the robustness of the estimates of the SFRs, stellar masses and dust attenuation parameters of 13 sources in the CEERS field observed with HST, JWST/NIRCam and NIRSpec between redshifts 1 and 3. These objects were selected based on the availability of at least one Balmer and one Paschen line with S/N 5 to allow for a solid determination of the amount of differential attenuation at play in our sample. The aforementioned physical parameters were computed using a newly-developed version of CIGALE which combines photometric fluxes and emission lines EWs as input data. We summarise our main results as follows:
The absorption line EWs of 7 hydrogen lines from Pa (up to 2 ) to H (up to 7 ) are measured. This is the first time an absorption EW for the Pa line is measured and a steady continuum of decreasing EWs with longer wavelengths is found.
The photometric data is better fitted with BC03 compared to BPASS while the quality is similar for EWs. Meanwhile, the SED fitting code overestimates the EWs when fitting photometry only (by an average of 15%), an effect we find to be even more pronounced when using the BPASS models (average of 45%).
We find higher SFRs (0.34 dex) and lower stellar masses (0.17 dex) with the fit of spectro-photometric data using BPASS compared to BC03. There is an overestimation of the SFRs (0.16 dex with BC03, 0.03 with BPASS) and an underestimation of stellar masses (0.15 dex) when fitting only photometry compared to fitting spectro-photometric data.
From the best model computed by CIGALE for each source, we assess the contribution of the H line emission to the measurements made with the NIRCam filters (F150W or F200W, depending on the redshift). We find H to be a major contributor to the broadband photometric fluxes ( average of 36%, up to 68%), even more so while using the BPASS models (average of 42%, up to 76%).
We find that using a double power law as introduced by the CF00 recipe to model the dust attenuation allows for solid constraints on the differential attenuation. The attenuation parameters derived for a subsample of sources with ¿ 0.3 mag are, on average, consistent with the CF00 recipe, especially the differential attenuation (=0.37 0.08). The change of flexibility in the dust attenuation recipe has no effect on SFRs and stellar masses, but this could be different for more attenuated objects. We reconstruct the effective attenuation curves of our subsample and derive a unique power law having an exponent of -0.96.
@@ Acknowledgements.
- Bagley et al. (2022) Bagley, M. B., Finkelstein, S. L., Koekemoer, A. M., et al. 2022, arXiv e-prints, arXiv:2211.02495
@@
- Bolamperti et al. (2023) Bolamperti, A., Zanella, A., Meštrić, U., et al. 2023, MNRAS[arXiv:2310.04498]
- Boquien et al. (2022) Boquien, M., Buat, V., Burgarella, D., et al. 2022, arXiv e-prints, arXiv:2202.11723
@@
- Finkelstein et al. (2022) Finkelstein, S. L., Bagley, M. B., Ferguson, H. C., et al. 2022, arXiv e-prints, arXiv:2211.05792
@@
- Osborne & Salim (2024) Osborne, C. & Salim, S. 2024, arXiv e-prints, arXiv:2401.06865
@@
- Robertson et al. (2024) Robertson, C., Holwerda, B., Young, J., et al. 2024, arXiv e-prints, arXiv:2403.15619

### v3

@@ 6.2 Derivation of the effective attenuation curves
With CIGALE, we can derive the effective attenuation curve for each galaxy of our subsample, being the total attenuation in the V band. We select fifteen wavelength bands (from the rest-frame far-UV to the near-IR) as the code is able to compute the attenuation in filters different from the ones used to fit the input data. With this feature in mind we add fifteen custom-made filters to the code. We choose to create very narrow filters (10 nm each) to carefully select which part of the spectrum is going to be sampled. The reason for this choice is two-fold: we can avoid contamination from emission and absorption lines but also pinpoint a specific part of the spectrum. We use the bayesian estimations of CIGALE for fifteen photometric bands to reconstruct the total effective attenuation curve of our subsample of galaxies.
@@
and we report the resulting exponents in Table 10. We also fit our whole subsample with a single powerlaw and find an exponent of =-0.96 0.10. By colour-coding each curve with its associated value of we see a trend between the decrease in and the steepening of the curve (i.e. the exponent is closer to -1.3). A high equals a negligible increase of attenuation for stars in birth clouds compared to the ISM meaning a single powerlaw (for ID 8588). The differential attenuation found for the galaxies with ¡ 0.35 mag leads to a steeper attenuation law. Using radiative transfer modelling, it has been shown that the effective attenuation curve depends on the amount of dust attenuation (e.g. Pierini et al., 2004; Saftly et al., 2015; Feldmann et al., 2017; Roebuck et al., 2019; Trayford et al., 2020). As predicted by the models the IR part of the curves do not follow the model of Calzetti et al. (2000) and ID 8588 is the only galaxy to follow the model in the UV-optical which shows the need for a two-components dust attenuation model. Finally we compared the values of obtained through a fit made with , , and being free and a fit made with and being free and both exponents fixed to -0.96. Both sets of values are consistent within their respective errorbars.
@@ 7 Impact of input data type and models on SFRs and stellar mass determinations
As getting robust SFRs and stellar masses is crucial in numerous contexts (e.g. Pacucci et al., 2023; Juodžbalis et al., 2024, submitted), we explore the effects of the choice of stellar populations to fit the spectro-photometric data on both aforementioned output parameters. Then we quantify the variation of SFRs and stellar masses between a fit performed with broad-band photometry alone and with spectro-photometric data, the latter being used as our reference fit. Finally we assess the impacts of the change of the attenuation recipe on the SFRs and stellar masses when fitting spectro-photometric data with the BC03 models.
@@ 7.1 Fits with BPASS or BC03 models
Here we compare SFRs and stellar mass estimations obtained with either BPASS or BC03 models. We start with SFRs estimations. Fig 12 shows a clear trend of increased SFRs when using BPASS compared to BC03 models. Excluding the three outliers (IDs 10293, 8736 and 5409 in red in Fig 12) the SFRs derived through SED fitting with BPASS are found on average 0.13 dex larger than with BC03. The value of the current SFR is expected to be sensitive to the burst fraction () added to the delayed SFH. The average value of with BPASS is found to be almost double the average value with BC03 (0.15 and 0.08 respectively) which may explain the higher SFRs we get while fitting our sample with the BPASS models. For the three outliers the difference is even higher as one of the two values is found to be null. The two galaxies with a much higher SFR obtained with BPASS are fitted with no burst with BC03 and with a value equal to 0.20 and 0.12 for ID 10293 and ID 8736 respectively with BPASS. Conversely for the galaxy with a much higher SFR with BC03 (ID 5409): no burst is found with BPASS while a small but positive burst fraction (0.01) is obtained with BC03. All in all the difference in SFRs obtained with both stellar population models seems to be correlated with the burstiness of the model rather than with an intrinsic difference between single and binary stars. Indeed Eldridge et al. (2017) and Wilkins et al. (2019) found that the SFRs obtained with BPASS are 0.1 dex only lower than those obtained with BC03 when the SFH is fixed.
We find higher stellar masses with BC03 models as shown in Fig 12. The mass derived through SED fitting with the BPASS models is on average 0.18 dex lower than with the BC03 models.Osborne & Salim (2024) also report lower masses (by 0.17 dex) when using the BPASS models compared to the BC03 models for a sample of CANDELS sources with redshift between 0.7 and 2.3. We attribute these lower masses to the deficit in low mass stars of the BPASS models addressed in Sect 4.2.
@@ 7.2 Fits of photometric data or spectro-photometric data
Here we assess the impact of using only photometry with both stellar population models on the stellar mass and SFR estimations. In Fig 13 we compare the estimates of the SFR obtained with a fit of photometric data to those from a fit of spectro-photometric data with both BC03 and BPASS. The SFR derived with the fit of photometric data only is on average 0.16 dex higher with BC03 and 0.03 dex higher with BPASS compared to values obtained with spectro-photometric measurements. Considering the scatter found in this comparison we also provide the maximal and minimal absolute value of the difference in . For BC03 (BPASS) the largest discrepancy is of 0.42 (0.60) dex and the lowest is of 0.04 (0.09) dex. Errors on the SFRs also become much lower when fitting the spectro-photometric data. SFRs lower than 30 M⊙/yr tend to be systematically overestimated when only photometric fluxes are fitted by 0.18 dex with BC03 and by 0.09 dex with BPASS. We attribute this increase to the stronger emission lines computed by the code in the absence of spectroscopic data as shown in Fig 9 which are due to a more active star formation process. Our results are in agreement with Pacifici et al. (2015) who also find a higher SFR (by 0.13 dex) while using BC03 and photometric data only for a sample of 364 3D-HST galaxies at z ¡ 3.
The stellar masses are found larger when EWs are added to the fit by 0.15 dex regardless of the chosen stellar populations as shown in Fig 13. The lower masses obtained with only photometric data are due to an overestimation of the line contribution (see Fig 9): the contribution of young stars to the SED is higher leading to a lower contribution of older stars which make the bulk of the galaxy mass. Our results are in agreement with the findings of Schaerer & de Barros (2010) who fits z 7 sources while Pacifici et al. (2015) finds no noticeable difference in the stellar mass output with the addition of spectroscopic data to the fits for their sample of z ¡ 3 galaxies.
@@ 7.3 Fits with different attenuation recipes
Dust attenuation has a direct impact on the derived SFRs and to a lesser extent on stellar masses. Any change in the shape of the attenuation curve can affect (albeit in varying degrees) the estimations of these quantities as shown in (e.g. Schaerer & de Barros, 2010; Pacifici et al., 2015; Lo Faro et al., 2017; Buat et al., 2018, 2019). In the following analysis we compare the effects of changing the attenuation curve parameters in the framework of the CF00 recipe.
We test three different attenuation curves with separate runs using the BC03 models only and the full spctro-photometric dataset. We do not include the BPASS models in these tests as the attenuation curves presented in Sect. 6.2 cannot be reconstructed with the binary stellar population models. The first run called is performed with the four parameters of the modified CF00 recipe taken free as presented in Table 4; it will be considered the reference run. The second run is performed with the original CF00 recipe () in which is fixed to 0.3 and both exponents are equal to -0.7 while the third one makes use of the single average power-law introduced in Sect. 6.2; is taken free and we call this run .
As already noted in Sect. 6.2 we do not find noticeable changes in between and . The SFRs are found consistent with a slightly larger scatter with CF00 than with the single power-law as shown in Fig 14. The agreement between the three stellar mass estimations is very good with a scatter of 0.03 and 0.06 dex for and respectively when they are compared to (see Fig 14).
From the results shown in this section we conclude that fixing the slope of the attenuation curve (and even ) is much less impactful on the SFRs and stellar masses than including binary stellar populations or fitting only photometric data. However this lesser effect could be a result of our sample being mostly composed of objects with low dust content. Indeed the most discrepant objects in Fig 14 have an amount of dust ¿ 1 mag.
@@ 8 Conclusions
In this work we assessed the robustness of the estimates of the SFRs, stellar masses and dust attenuation parameters of 13 sources in the CEERS field observed with HST, JWST/NIRCam and NIRSpec between redshifts 1 and 3. These objects are selected based on the availability of at least one Balmer and one Paschen line with S/N 5 to allow for a solid determination of the amount of differential attenuation at play in our sample. The aforementioned physical parameters are computed using a newly developed version of CIGALE which combines photometric fluxes and emission line EWs as input data. We summarise our main results as follows:
The absorption line EWs of seven hydrogen lines from Pa (up to 2 ) to H (up to 7 ) are measured. This is the first time an absorption EW for the Pa line is measured and a steady trend of decreasing EWs with longer wavelengths is found.
The photometric data is better fitted with BC03 compared to BPASS while the quality is similar for EWs. Meanwhile the SED fitting code overestimates the EWs when fitting photometry only (by an average of 15%) an effect we find to be even more pronounced when using the BPASS models (average of 45%).
We find higher SFRs (0.13 dex) and lower stellar masses (0.17 dex) with the fit of spectro-photometric data using BPASS compared to BC03. There is an overestimation of the SFRs (0.16 dex with BC03 and 0.03 with BPASS) and an underestimation of stellar masses (0.15 dex) when fitting only photometry compared to fitting spectro-photometric data.
From the best model computed by CIGALE for each source we assess the contribution of the H line emission to the measurements made with the NIRCam filters (F150W or F200W depending on the redshift). We find H to be a major contributor to the broadband photometric fluxes (average of 36% and up to 68%) even more so while using the BPASS models (average of 42% and up to 76%).
We find that using a double power law as introduced by the CF00 recipe to model the dust attenuation allows for solid constraints on the differential attenuation. The attenuation parameters derived for a subsample of sources with ¿ 0.3 mag are on average consistent with the CF00 recipe, especially the differential attenuation (=0.37 0.08). The change of flexibility in the dust attenuation recipe has no effect on SFRs and stellar masses but this could be different for more attenuated objects. We reconstruct the effective attenuation curves of our subsample and derive a unique power law having an exponent of -0.96.
@@ Acknowledgements.
- Bagley et al. (2023) Bagley, M. B., Finkelstein, S. L., Koekemoer, A. M., et al. 2023, ApJ, 946, L12
@@
- Bolamperti et al. (2023) Bolamperti, A., Zanella, A., Meštrić, U., et al. 2023, MNRAS, 526, 5263
- Boquien et al. (2022) Boquien, M., Buat, V., Burgarella, D., et al. 2022, A&A, 663, A50
@@
- Fernández et al. (2024) Fernández, V., Amorín, R., Firpo, V., & Morisset, C. 2024, arXiv e-prints, arXiv:2405.15072
@@
- Finkelstein et al. (2023) Finkelstein, S. L., Bagley, M. B., Ferguson, H. C., et al. 2023, ApJ, 946, L13
@@
- Osborne & Salim (2024) Osborne, C. & Salim, S. 2024, ApJ, 962, 59
@@
- Osterbrock & Ferland (2006) Osterbrock, D. E. & Ferland, G. J. 2006, Astrophysics of gaseous nebulae and active galactic nuclei
@@
- Robertson et al. (2024) Robertson, C., Holwerda, B. W., Young, J., et al. 2024, AJ, 167, 263